answer the problems using GRESA method:


1. A 53kg skateboarder pushes on the ground with -351N of force to set himself moving. How fast will the skateboarder accelerate?


2. A responsible driver, driving at a rate of 15m/s, sees the traffic light red. She presses on the brakes and after 3.20s, comes to a stop. The mass is 5400 g. How much force was exerted on the car?


1
Expert's answer
2021-03-14T19:14:28-0400

1) Given:

"m=53\ kg, F=351\ N."

Required:

Acceleration of the skateboarder.

Equation:

"F=ma."

Solution:

"a=\dfrac{F}{m}=\dfrac{351\ N}{53\ kg}=6.62\ \dfrac{m}{s^2}."

Answer:

"a=6.62\ \dfrac{m}{s^2}."


2) Given:

"v_i=3.2\ \dfrac{m}{s}, v_f=0, t=3.2\ s, m=5.4\ kg."

Required:

Force exerted on the car.

Equation:

"F\Delta t=m(v_f-v_i)."

Solution:

"F=\dfrac{m(v_f-v_i)}{\Delta t},""F=\dfrac{5.4\ kg\cdot(0-15\ \dfrac{m}{s})}{3.2\ s}=-25.31\ N."

Answer:

"F=-25.31\ N."